# MROSS: Multi-Round Region-Based Optimization for Scene Sketching

### v1

@@ Introduction
Figure 9: Comparison of different loss function. The effect of using only the CLIP-Based Semantic loss: Capture abstract semantic details, like background, light and shadow. The effect of using only the VGG-Based Feature loss: emphasize image features, such as geometric contours.
Ours v.s. CLIPasso Ours v.s. Photo-Sketching Ours v.s. Chan et al. Ours v.s. UPDG
Ours 93.0% 96.6% 42.7% 29.8%
Others 2.7% 3.4% 32.9% 24.1%
Equal 4.3% 0.0% 24.4% 46.1%
Table 2: Results of our user study. For each method, we specify the percent of responses that preferred our sketch, the sketch of the alternative method, or found the sketches to be similar in their ability to capture the scene semantics.
Quantitative Evaluation
LPIPS. LPIPS (Zhang et al. 2018) measures the perceptual similarity between two images and lower LPIPS scores indicate less perceptual difference. We calculate the scores between input images and sketches of different methods. Table 1 shows the average LPIPS scores for each methods. Our results under high level of abstraction are close to those of Chan et al. (Chan, Durand, and Isola 2022), which means our sketches follow the content of the input image. The results of XoG (Winnemöller, Kyprianidis, and Olsen 2012) are not the best, it may because LPIPS does not only consider edge information.
SSIM. To measure the fidelity level of sketches, we compute the SSIM (Wang et al. 2004) score between each input image and the corresponding sketch. In Table 1, we show the average resulting scores, where a higher score indicates a higher similarity. The sketches by XoG (Winnemöller, Kyprianidis, and Olsen 2012) obtained high scores, which is consistent with our observation that SSIM highlight the edges of the image. Under different levels of abstraction, our method shows relatively good scores, indicating that we have a better effect on edge information preservation. This shows the effectiveness of our edge-based sampling method.
User Study. The user study examines how well the sketches depict the input scene. Since the absence of CLIPascene’s (Vinker et al. 2022a) code, we use 30 scene images to compare our sketches with four methods: CLIPasso (Vinker et al. 2022b), Photo-Sketching (Li et al. 2019), Chan et al. (Chan, Durand, and Isola 2022) and UPDG (Yi et al. 2020).
@@
The participants were presented with the input image along with two sketches, one produced by our method and the other by the alternative method. In order to make a fair study, we compared Photo-Sketching (Li et al. 2019) with our sketches at the lower abstraction level. Conversely, we compared the sketches of Chan et al. (Chan, Durand, and Isola 2022) and UDPG (Yi et al. 2020), more detailed. Table 2 shows the average resulting scores among all participants. Compared to CLIPasso (Vinker et al. 2022b) and Photo-Sketching (Li et al. 2019), our method achieves significantly higher rates. As we can see, although sketches produced by Chan et al. (Chan, Durand, and Isola 2022) and UPDG (Yi et al. 2020) are highly detailed, 42.7% responses preferred our sketches when compared to Chan et al. (Chan, Durand, and Isola 2022) and 46.1% of responses considered our sketches is similar to UPDG’s (Yi et al. 2020). The results of user study demonstrate that sketches produced by our method well represent the elements of the input scene.
Ablation Study
In Figure 9, we show the results of different loss function. As can be seen, in that case the VGG-Based Feature loss faithfully preserves geometric structure, while the CLIP-Based Semantic loss strongly conveys the semantic concept, however at the cost of structure. Using their combination can result in an ideal sketch, which is not only semantically accurate, but also retains geometric details.
Conclusion
@@
We propose a multi-round optimization for scene sketching based on different regions. We progressively infers the sketch with the help of proposed stroke initialization and loss functions: a CLIP-Based Semantic loss and a VGG-Based Feature loss. We can investigate the coarse-to-concrete representation for complex images with different regions. Extensive experiments and human evaluation confirm the superior reconstruction performance of our method over the existing. In the emerging field of sketch generation, we hope our work will open the door for further research.
Acknowledgments
@@
This research was supported by National Natural Science Foundation of China No.62202151 and No.62202152.

### v2

@@ III-C Comparison to State-of-the-art Methods
Figure 8: Ablation of loss function. The effect of using only the CLIP-based Semantic Loss: capture abstract semantic details, like background, light and shadow. The effect of using only the VGG-based Feature Loss: emphasize structural features.
III-D User Study and Results
@@
User study. To evaluate how well the sketches depict the input scene, we conducted a user study. We use 30 scene images to compare our sketches with six methods: CLIPasso [8], CLIPascene [14], XoG [21], Photo-Sketching [22], Chan et al. [23] and UPDG [24]. The participants were presented with the input image along with two sketches, one produced by our method and the other by the alternative method. In order to make a fair study, we compared Photo-Sketching with our sketches at the lower abstraction level. Table II presents the final preference rates, showing MROSS gains higher acceptance from participants.
Results. Fig. 6 and Fig. 7 demonstrate comparisons of MROSS with prior state-of-the-art methods. In Fig. 6 we select XoG [21], Photo-Sketching [22], Chan et al. [23], and UPDG [24], which do not have the ability to adjust the sketch abstraction level or generate vector-based sketches, for comparison. Given the different levels of abstraction of these methods, we choose to demonstrate the different abstraction level generation of our method by generating three different number of strokes (128, 64 and 32). In Fig. 7, we provide comparisons with vector-based methods CLIPasso [8] and CLIPascene [14]. Both can generate sketches at different abstraction levels and generate sketches in vector format. Therefore, we set the number of strokes at different abstraction levels (128 and 32) among these two methods for better comparison.
In contrast, MROSS, benefiting from alignment in structure and sematics, is able to maintain recognizability, underlying structure, and essential visual components of the scene drawn regardless of the level of abstraction. CLIPasso [8] aims to portray objects accurately. However, for the scene, it fails to depict the total image in details. Then we compared to CLIPascene [14], for a fair comparison, we use the same decomposition technique to separate the input images into foreground and background. Then we use our method to sketch each part separately before combining them. CLIPascene [14] captures the whole image, which prevents them from emphasizing important regions of the input image. As shown in Fig. 7, our method is able to capture the entire image while displaying contours and details of key objects.
III-E Ablation Study
In Fig. 8 we present a comprehensive analysis of the results obtained by the results of various loss functions. Fig. 8 clearly shows that CLIP-based Semantic Loss can robustly capture semantic concepts but falls short in terms of structural fidelity. VGG-based Feature Loss could perfectly preserve the geometric structure, ensuring a faithful representation of the original form. Ablating these loss functions can prove that our final loss combination not only achieves semantic accuracy but also effectively preserves complex geometric details, achieving a harmonious balance between semantic expression and structural integrity.
IV CONCLUSION
@@
We propose MROSS, an innovative scene sketching method that adopts a multi-round optimization strategy focusing on autonomously selected different regions. This enables a coarse-to-concrete representation, which is particularly beneficial for processing complex images. Meanwhile, we guide sketch generation with CLIP-based Semantic Loss and VGG-based Feature Loss. Extensive experiments and user studies confirm the superior scene generation performance of MROSS compared to existing methods.
V ACKNOWLEDGEMENT
@@
This research was supported by National Natural Science Foundation of China (Grant No., 62202152 & 62202151).